Question 1 Write each value as a reduced fraction, decimal value, and a percentage term 0.02 5%
Answer #1

1) 1/4 = 0.25 = 25%

Therefore, Decimal value is 0.25 and percentage term is 25%.

2) 0.02 = 2/100 = 1/50 = 2%

Reduced fraction is 1/50. Percentage term is 2% and decimal value is already given.

3) 5% = 5/100 = 1/20 = 0.05

Reduced fraction is 1/20 and decimal value is 0.05.